a meter stick of mass 400g is pivoted at one end displaced through an angle of 60 degree .then what amount of potential energy increases

Grade:11

2 Answers

Vikas TU
7 years ago
Initial P.E = mgl/2 = > 0.4*10*1/2 = 2 Joules.
At 60 degree angle ,
P.E. = mgl(1-cos60)/2 = > 0.4*10*1(1-0.5)/2  => 1 Joules.
Hence the potential energy get decreased by = 2 – 1 = 1 Joue instead of increasing.
ankit singh
3 years ago

 

Here 
Mass of stick ,m=0.4 kg
Length of stick ,  l=1m
Initial Potential energy =2mgl=20.4×10×1=2 J
 
At 600 angle ,
 
Final Potential energy  =2mgl(1cos600)=20.4×10×1(10.5)=1 J.
 
So , Its potential energy decreased by 21=1 J